Let X1 and X2 be any two random variables, then E( Cov( X1, X2) ) =

A.

X1 times X2, i.e., X1*X2.

B.

X1 / X2.

C.

Cov( X1, X2).

D.

none of the above
Answer #1

The answer is:

{\color{DarkRed} \text{ C. Cov( X1, X2).}}

Proof:

Cov( X_1, X_2) = E(X_1,X_2) - E(X_1)*E(X_2)

\rightarrow E(Cov( X_1, X_2)) = E(E(X_1,X_2) - E(X_1)*E(X_2))

\rightarrow E(Cov( X_1, X_2)) = E(X_1,X_2) - E(E(X_1))*E(E(X_2))

\rightarrow E(Cov( X_1, X_2)) = E(X_1,X_2) -E(X_1)*E(X_2)

\rightarrow E(Cov( X_1, X_2)) = Cov( X_1, X_2)

Let me know in the comments if anything is not clear. I will reply ASAP! Please do upvote if satisfied!